OpenStudy (anonymous):

Find the determinant of the matrix [-4 -4] [ 5 -6]

OpenStudy (the_fizicx99):

 \(\ (-4)(-6) = 24 | (-4)(5) = -20 \) \(\ 24 - (-20) = 24 + 20 = 44 \)